# Optimal Pure Differentially Private Sparse Histograms in Near-Linear Deterministic Time

### v1

@@ 4.2.2 Utility Guarantee
Recall from Theorem 5.2, for each $t\in\mathbb{N}_{\leq{n}}$ and each $\beta^{\prime}>2\gamma^{\prime}\doteq 2\varepsilon\gamma/d$, $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\gamma^{\prime}})}}\left({t}\right)$ is $(\alpha,\beta)$-accurate estimator of $t$ for $\alpha\doteq{\lceil{\frac{1}{\varepsilon}\cdot\ln\frac{4}{\beta^{\prime}}}\rceil}\in O{({\frac{1}{\varepsilon}\cdot\ln\frac{1}{\beta^{\prime}}})}$.
@@
Now fix any $\beta>2\varepsilon\gamma$, and let $\beta^{\prime}\doteq\beta/d>2\gamma^{\prime}$. Then $\alpha\in O{({\frac{1}{\varepsilon}\cdot\ln\frac{d}{\beta}})}$. We consider the accuracy of each coordinate $\tilde{h}[i]$ across three cases.
@@
Case 1 ($i\in I$): In this case, $\tilde{h}[i]$ is directly constructed from the output of $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\gamma^{\prime}})}}{({\vec{h}[i]})}$, and is thus an $(\alpha,\beta^{\prime})$-accurate estimator of $\vec{h}[i]$.
Case 2 ($i\in[d]\setminus{({\operatorname{\textsc{supp}}{({\vec{h}})}\cup I})}$): Here, $i$ is neither in the support of $\vec{h}$ nor in $I$, so $\tilde{h}[i]=0=\vec{h}[i]$, and the error is exactly zero.
@@
Case 3 ($i\in\operatorname{\textsc{supp}}{({\vec{h}})}\setminus I$): In this case, $\hat{h}[i]$ is an $(\alpha,\beta^{\prime})$-accurate estimator of $\vec{h}[i]$. Moreover, the fact that $i\notin I$ implies that $\hat{h}[i]<\tau$. Hence,
@@
$\begin{array}[]{c}\lvert{\vec{h}{\left[{i}\right]}-\tilde{h}{\left[{i}\right]}}\rvert=\vec{h}{\left[{i}\right]}\leq\lvert{\vec{h}{\left[{i}\right]}-\hat{h}{\left[{i}\right]}}\rvert+\hat{h}{\left[{i}\right]}=\alpha+\tau\in O{({\frac{1}{\varepsilon}\cdot\ln\frac{d}{\beta}})}.\end{array}$
@@ 4.2.3 Running Time
First, as shown in Theorem 5.2, the noise sampler $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}$ has initialization time $\tilde{O}(1)$ and uses $O\bigl{(}(\tfrac{1}{\varepsilon}+\ln\tfrac{d}{\varepsilon\gamma}+\ln n)\cdot(\ln\tfrac{1}{\varepsilon}+\ln\tfrac{d}{\varepsilon\gamma}+\ln n)\bigr{)}$ bits of memory.
Under the assumption in Theorem 4.1 that the descriptions of $n$, $\varepsilon$, and $\gamma$ fit in a constant number of machine words (i.e., $\ln n,\ln\tfrac{1}{\varepsilon},\ln\tfrac{1}{\gamma}\in O(\omega)$, where $\omega$ is the word size), the total memory usage is $O\bigl{(}\tfrac{1}{\varepsilon}+\ln\tfrac{d}{\varepsilon\gamma}+\ln n\bigr{)}$ words.
After initialization, each sample from the noise distribution can be generated in $O\bigl{(}\tfrac{1}{\omega}\cdot(\ln\tfrac{1}{\varepsilon}+\ln\tfrac{d}{\varepsilon\gamma}+\ln n)\bigr{)}=O(1)$ time. Hence, generating $n+k$ such samples requires total time $O(n+k)$.
@@
To sample a uniform random subset from $\binom{[d]\setminus I_{1}}{n+k-\lvert{I_{1}}\rvert}$ (the set of all subsets of size $n+k-\lvert{I_{1}}\rvert$ from $[d]\setminus I_{1}$), we first sample a uniform random subset of size $n+k$ from $[d]$, and then randomly select $n+k-\lvert{I_{1}}\rvert$ elements from this subset that do not appear in $I_{1}$.
@@
    Abort and output a fixed sparse histogram. Terminate Algorithm 3 early and return a fixed sparse histogram, e.g., one where elements $1$ through $n$ each have count $1$. Although it may slightly degrade utility, it does not compromise privacy. We show that this increases the failure probability of the utility guarantee by at most an additive $\sqrt{n+k}\cdot\exp(-(n+k)/16)$.
@@
To analyze this, we apply the Poisson approximation method (Mitzenmacher and Upfal, 2005). Let $Y_{1},\ldots,Y_{d}$ be independent Poisson random variables with mean $m/d$, i.e., $\Pr\left[Y_{i}=t\right]=e^{-m/d}\cdot\frac{(m/d)^{t}}{t!}$ for each $t\in\mathbb{N}$. Define $S_{Y}=\sum_{i\in[d]}\mathds{1}_{\left[Y_{i}>0\right]}$ analogously. The Poisson approximation technique implies that: $\Pr\left[S_{Y}\leq n+k\right]\leq\sqrt{m}\cdot\Pr\left[S_{Y}\leq n+k\right]$. Therefore, it suffices to bound $\Pr\left[S_{Y}\leq n+k\right]$.
Since $\Pr\left[Y_{i}>0\right]=1-e^{-\frac{m}{d}}$, we have $\operatorname*{\mathbb{E}}\left[S_{Y}\right]=d\cdot{({1-e^{-\frac{m}{d}}})}\geq d\cdot\frac{m}{2d}=m/2$, where the inequality holds whenever $m\leq d$. Via Chernoff bound,

### v2

@@ 4.2.2. Utility Guarantee
Consider any $\beta>2\varepsilon\gamma$. Set $\gamma^{\prime}\doteq\varepsilon\gamma/d$ and $\beta^{\prime}\doteq\beta/d$, so that $\beta^{\prime}>2\gamma^{\prime}$. Recall from Theorem˜1.4 that for each $t\in\mathbb{N}_{\leq{n}}$, $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\gamma^{\prime}})}}\left({t}\right)$ is an $(\alpha,\beta^{\prime})$-accurate estimator of $t$, where
$\alpha\in O{\left({\frac{1}{\varepsilon}\ln\frac{1}{\beta^{\prime}}}\right)}=O{\left({\frac{1}{\varepsilon}\ln\frac{d}{\beta}}\right)}.$
@@
We consider the accuracy of each coordinate $\tilde{h}[i]$ across three cases.
Case 1 ($i\in I$)
@@
: In this case, $\tilde{h}[i]$ is directly constructed from the output of $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\gamma^{\prime}})}}{({\vec{h}[i]})}$, and is thus an $(\alpha,\beta^{\prime})$-accurate estimator of $\vec{h}[i]$.
Case 2 ($i\in[d]\setminus{({\operatorname{\textsc{supp}}{({\vec{h}})}\cup I})}$)
: Here, $i$ is neither in the support of $\vec{h}$ nor in $I$, so $\tilde{h}[i]=0=\vec{h}[i]$, and the error is exactly zero.
Case 3 ($i\in\operatorname{\textsc{supp}}{({\vec{h}})}\setminus I$)
@@
: In this case, $\hat{h}[i]$ is an $(\alpha,\beta^{\prime})$-accurate estimator of $\vec{h}[i]$. Moreover, the fact that $i\notin I$ implies that $\hat{h}[i]<\tau$. Hence,
@@
$\lvert{\vec{h}{\left[{i}\right]}-\tilde{h}{\left[{i}\right]}}\rvert$ $=\vec{h}{\left[{i}\right]}\leq\lvert{\vec{h}{\left[{i}\right]}-\hat{h}{\left[{i}\right]}}\rvert+\hat{h}{\left[{i}\right]}=\alpha+\tau\in O{\left({\frac{1}{\varepsilon}\cdot\ln\frac{d}{\beta}}\right)}.$
@@ 4.2.3. Running Time
Under the assumption in Theorem˜4.1 that the descriptions of $n$, $\varepsilon$, and $\gamma$ fit in a constant number of machine words (i.e., $\ln n,\ln\tfrac{1}{\varepsilon},\ln\tfrac{1}{\gamma}\in O(\omega)$, where $\omega$ is the word size), Theorem 1.4 implies that the noise sampler $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}$ has initialization time $\tilde{O}(1)$ and uses $O\!\left(\tfrac{1}{\varepsilon}+\ln\tfrac{d}{\varepsilon\gamma}+\ln n\right)$ words of memory.
After initialization, each sample from the noise distribution can be generated in $O(1)$ time. Thus, producing $n+k$ samples takes total time $O(n+k)$.
@@
To sample a uniform random subset from $\binom{[d]\setminus I_{1}}{n+k-\lvert{I_{1}}\rvert}$ (the set of all subsets of size $n+k-\lvert{I_{1}}\rvert$ from $d]\setminus I_{1}$), we first sample a uniform random subset of size $n+k$ from $[d]$, and then randomly select $n+k-\lvert{I_{1}}\rvert$ elements from this subset that do not appear in $I_{1}$.
@@
    Abort and output a fixed sparse histogram. Terminate Algorithm˜2 early and return a fixed sparse histogram, e.g., one where elements $1$ through $n$ each have count $1$. Although this may slightly degrade utility, it does not compromise privacy, since sampling $m$ independent elements from $[d]$ is independent of the input dataset. We show that this increases the failure probability of the utility guarantee by at most an additive $\sqrt{n+k}\cdot\exp(-(n+k)/16)$.
@@
To analyze this, we apply the Poisson approximation method (Mitzenmacher and Upfal, 2005). Let $Y_{1},\ldots,Y_{d}$ be independent Poisson random variables (Definition˜A.8) with mean $m/d$, i.e., $\Pr\left[Y_{i}=t\right]=e^{-m/d}\cdot\frac{(m/d)^{t}}{t!},\quad\forall t\in\mathbb{N}.$ Define $S_{Y}=\sum_{i\in[d]}\mathds{1}_{\left[Y_{i}>0\right]}$ analogously. The Poisson approximation theorem (Theorem˜A.9) implies that
$\Pr\left[S_{X}\leq n+k\right]\leq e\sqrt{m}\cdot\Pr\left[S_{Y}\leq n+k\right].$
Thus, it suffices to bound $\Pr\left[S_{Y}\leq n+k\right]$. Since $\Pr\left[Y_{i}>0\right]=1-e^{-m/d}$, we have $\operatorname*{\mathbb{E}}\left[S_{Y}\right]=d\cdot{({1-e^{-m/d}})}\geq d\cdot\frac{m}{2d}=m/2,$ where the inequality uses $1-e^{-x}\geq x/2$ for $0\leq x\leq 1$, applicable whenever $m\leq d$.
Via the Poisson Chernoff bound (Theorem˜A.10),